# DCT-MARL: A Dynamic Communication Topology Based MARL Algorithm for Platoon Control

### v1

@@ IV DCT-MARL Based Algorithm
In this section, the proposed DCT-MARL algorithm is described, as shown in Figure.1. The platoon control task is then formulated as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP). Furthermore, a dynamic communication topology mechanism is incorporated into the DCT-MARL framework to enhance the robustness of control performance under non-ideal communication conditions.
@@
The proposed DCT-MARL algorithm is built upon the actor-critic architecture. Each actor consists of a communication network, which is a multi-key gated network based on an inference mechanism. and an action policy network, which determines control actions. Each agent acquires state information by communicating with vehicles within its communication range. Considering potential communication failures, each agent dynamically selects a communication topology at every control cycle based on its local observation, integrating its own state to make control decisions.
@@
The centralized critic receives the joint actions and state representations from all agents as input to estimate each agent’s action-value function at each time step. An attention mechanism is utilized to extract relevant information from other agents’ observation-action pairs, enabling each agent to selectively filter critical environmental information. The critic network is trained via temporal difference learning[34] and is only used during training. This paper optimize agent $i$’s policy by minimizing the Kullback–Leibler (KL) divergence $D_{KL}(q_{\theta_{i}}\|p)$, which is given by:
$-D_{\text{KL}}(q_{\theta_{i}}\parallel p)=$ $\sum_{t}\mathbb{E}_{s_{t},u_{t},A_{t}\sim q_{\theta_{i}}}\Big{[}r_{i}(s_{t},u_{t})+$ (8)
$H\big{(}\pi_{i}(u_{i}\mid o_{i}(s,A))\cdot\rho(A_{t}\mid s_{t})\big{)}\Big{]}.$
Beyond the reward term, the objective introduces an additional conditional entropy term for the joint distribution $\pi_{i}(a_{i}|o(s_{t},A_{t}))\rho(A_{t}|s_{t})$, which encourages exploration for agent $i$ and enhances graph policy $\rho$. The objective function for agent $i$ is now redefined as:
@@
The loss function for $Q_{i}$ is defined as:
@@
During training, we follow these iterative procedures:
@@
where, $y_{i}=\begin{cases}1,&\text{if }P_{\omega}(y_{i}\mid o_{i})\geq\delta\\ 0,&\text{otherwise},\end{cases}$ and $\delta$ is a hyperparameter. Notably, if a particular agent plays a critical role in shaping agent $i$’s state estimation during centralized training, its information should remain accessible during decentralized execution. To this end, a dynamic communication topology based on communication target selection is proposed and shared across both training and execution phases, enabling robust and adaptive coordination.
The proposed mechanism employs a multi-key gating structure to dynamically select communication targets, accommodating practical V2V constraints such as delay and packet loss. By adjusting the topology in response to traffic context, it ensures stable coordination and safety performance of platoon systems under complex and dynamic environments.
@@ IV-A Dec-POMDP Formulation
State Space $S$: To mitigate the performance degradation caused by communication delay, this paper extend the definition of the agent’s state space by incorporating both historical control information and delay descriptors. Specifically, at time step $t$, the state of vehicle $i$ is defined as:
$S_{i,t}=\left\{p_{i,t},v_{i,t},u_{i,t},u_{t-1},\xi_{i,t}\right\},$ (14)
where $p_{i,t}$, $v_{i,t}$, and $u_{i,t}$ denote the vehicle’s current position, velocity, and acceleration, respectively; $u_{t-1}$ represents the most recent control input, and $\xi_{i,t}$ denotes the communication delay associated with the latest received information.
Although this formulation deviates from the conventional state definition, it is a rational extension under communication-impaired environments, where both the delayed control input and latency should be explicitly considered in the state representation. In connected vehicle systems, time-varying wireless communication often causes neighboring information to arrive with variable delay. If the control policy relies solely on such delayed observations, the Markov property may be compromised, thereby impairing decision-making performance.
@@
Observation Space $O$: Agents use communication to obtain information beyond their direct observations to assist decision-making. Each agent’s local observation $o_{i}\in S_{i}$ is a partial description of the environment state. At time step $t$, each agent $j$ in the platoon transmits its driving information $X_{j,t}={p_{j,t},v_{j,t},u_{j,t},t_{j,t}}$, neighboring agent $i$ obtains driving information from multiple vehicles within its V2V communication range, forming its partial observation $o_{i}$ and determining which agents are within communication range. This solves the problem of possible communication failures. For example, if agent l fails within the current control interval k, agent i cannot observe agent l’s information through V2V communication, and can receive the status information of other related vehicles for control output. It avoids the problem that the amount of states received may be insufficient due to communication failures. The global state $O$ is then defined as the combination of all agents’ local observations: $O=\{o_{1},o_{2},\dots,o_{n}\}.$ Similarly, the global state $S$ is defined as: $S=\{s_{1},s_{2},\dots,s_{n}\}.$
@@
$r_{i,t}=-\big{\{}$ $\omega_{1}\left(\text{err}_{p_{i},t}\right)+\omega_{2}\left(\text{err}_{v_{i},t}\right)+\omega_{3}\left(\text{jerk}_{i,t}\right)$ (15)
$+\omega_{4}\cdot\operatorname{ReLU}\left(d_{\text{safe}}-d_{i,t}\right)\big{\}},$ (16)
@@
$\text{jerk}_{i,t}=\frac{acc_{i}(t)-acc_{i}(t-1)}{\Delta t}.$ (17)
@@
In addition, considering the communication delay problem, each agent can only access the lagged observation information when making a decision, and cannot directly obtain the current true state. Therefore, we modify the reward function based on conditional expectation to adapt it to the optimal learning in the communication constrained environment. Specifically, for each agent $i$, in the extended state $S_{i,t}$, the reward function of delay is defined as follows:
$\tilde{R}_{i,t}\left(S_{i,t},u_{i,k}^{\mathrm{CL}}\right)=\mathbb{E}_{x_{i,t}}\left[R_{i,t}\left(x_{i,t},u_{i,t}\right)\mid S_{i,t}\right].$ (18)
@@
IV-B Dec-POMDP with Dynamic Directed Graph
@@
Based on Dec-POMDP, we propose a communication topology using dynamic directed graphs. The structural properties of the graph are represented by the adjacency trajectory matrix $A$, which is denoted as $\mathbf{A}\in\mathbb{B}^{l\times n\times n}$, where B indicates the Boolean matrix, l denotes the length of the trajectory and n indicates the number of agents. If $\lceil A^{ij}_{t}\rceil=0$, then no communication occurred from agent $i$ to agent $j$ at time $t$. If $\lceil A^{ij}_{t}\rceil=1$, then communication transpired from agent $i$ to agent $j$ at time $t$.
IV-C Multi-Key Gating Network Based Dynamic Communication
@@
During communication, an agent intuitively prefers to interact with those agents that significantly influence its decision-making strategy. Specifically, assuming agent $i$ can receive information from agent $j$, it first inputs its local observation $o_{i}$ along with the ID $d_{j}$ of agent $j$ into the model, which then outputs a confidence score indicating the likelihood of communication with agent $j$. To this end, this paper employs a feedforward neural network to implement causal inference, leveraging the KL divergence to quantify the causal confidence between agent $i$ and other agents. The formulation is given as follows:
$D_{\text{KL}}(P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})\|P(a_{i},a_{j}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})),$ (20)
@@
where $P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ and $P(a_{i},a_{j}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ denote two conditional probability distributions over the action space of agent $i$, and $\boldsymbol{o}$ represents the joint observations of all agents. Here, $\boldsymbol{a}_{-i}$ refers to the joint actions of all agents excluding agent $i$, and similarly, $\boldsymbol{a}_{-ij}$ denotes the joint actions of all agents excluding both agents $i$ and $j$.
@@
The latter probability distribution does not depend on agent $j$’s actions, meaning that when making decisions, agent $i$ ignores the actions of agent $j$. Note that the conditional actions of other agents $a_{-i}$ are sampled from their current policies. The magnitude of KL divergence quantifies how agent $i$ would adjust its own policy if it considered agent $j$’s actions, thus indicating the degree of correlation between agent $j$’s policy and agent $i$’s policy.
@@
The two probability distributions are computed using the joint action-value function, where $P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ is derived from the softmax distribution over agent $i$’s actions:
$P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})=\frac{\exp(Q(a_{i},\boldsymbol{a}_{-ij},\boldsymbol{o})/\lambda)}{\sum_{a_{i}^{\prime}}\exp(Q(a_{i}^{\prime},\boldsymbol{a}_{-ij},\boldsymbol{o})/\lambda)},$ (21)
@@
where $Q(\boldsymbol{a},\boldsymbol{o})$ is the joint action-value function, and $\lambda\in\mathbb{R}^{+}$ is the temperature parameter. The marginal distribution $P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ is computed as follows:
$P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})=\sum_{a_{j}}P(a_{i},a_{j}\mid\boldsymbol{a}_{-ij},\boldsymbol{o}).$ (22)
@@
Then, a multi-key gating network is used to select $m$ communication targets, where the Gumbel-Softmax activation function determines whether an agent should communicate or not. The computation is given by:
$key_{k}^{i}=\text{ gumbel }-\operatorname{softmax}\left(\mathrm{W}_{\mathrm{v}}\tanh\left(\mathrm{~W}_{\mathrm{q}}\mathrm{o}_{\mathrm{k}-1}^{\mathrm{i}}+\mathrm{W}_{\mathrm{k}}\mathrm{o}_{\mathrm{k}-1}^{\mathrm{i}}\right)\right),$ (23)
where Gumbel-Softmax [35, 36]is an activation function with two output dimensions (communicate or not). The weight matrices $\mathbf{W}_{v}$, $\mathbf{W}_{q}$, and $\mathbf{W}_{k}$ determine the communication importance, considering the dynamically changing communication links in a vehicle platoon scenario due to traffic conditions.
Finally, the adjacency matrix $\mathbf{A}$ is generated by summing over all keys:
$A_{i,k}=\left\lceil\max\left(key_{k}^{i}\mid j\right)\right\rceil,$ (24)
where $key_{k}^{i}\mid j$ represents the total number of keys obtained by agent $i$ in control cycle $k$, and $\lceil\cdot\rceil$ denotes the ceiling function. This ensures that the final communication topology is dynamically adjusted rather than fixed. Additionally, the number of keys ensures that each agent only communicates with a limited number of selected agents, thereby maintaining a stable input dimension.
@@
To address the lack of interpretability in purely data-driven methods, this paper introduce a dynamic weighting module based on safety distance, which assigns higher weights to closer vehicles to reflect their critical role in collision avoidance. Specifically, the weight of vehicle $j$ is defined as:
$\omega_{j,k}=\frac{\frac{1}{2^{|i-j|}}}{\sum_{l\in N_{\mathrm{i},\mathrm{k}}}\frac{1}{2^{|\mathrm{i}-\mathrm{j}|}}},$ (25)
where $d_{ij}$ represents the distance between vehicle $i$ and vehicle $j$, and $\epsilon$ is a small positive constant to avoid division by zero. Finally, all received messages are fed into the message encoder network $e_{i}(\cdot)$ to generate encoded messages $c_{i}$, and the policy network produces the action distribution $\pi_{i}(a_{i}|c_{i},o_{i})$.
@@ V-A Experimental Setup
To validate the effectiveness of the proposed control strategy, this paper further incorporates empirical trajectory data from the NGSIM dataset[37], a widely used benchmark in vehicle platoon control research that includes congested traffic scenarios. This enables the evaluation of the controller’s ability to handle stop-and-go traffic oscillations commonly observed in real-world platoon environments. Specifically, the driving profile of the leading vehicle is extracted from real-world trajectory data collected on the eastbound I-80 highway in the San Francisco Bay Area. The use of realistic driving patterns contributes to more effective policy learning for connected vehicles platoon operating under complex and dynamic conditions.
In the communication simulation, the wireless channel model incorporates both small-scale and large-scale fading effects. Small-scale fading is modeled as Rayleigh fading, capturing rapid signal fluctuations due to multipath propagation. Large-scale fading includes path loss (PL) and shadow fading (SF).
Path loss is modeled using the WINNER+ B1 model recommended by 3GPP [38], which defines as:
$\mathrm{PL}(d)=A_{0}+10n\log_{10}\left(\frac{d}{d_{0}}\right)+X_{\sigma},\quad\text{(dB)}$ (26)
where $d_{0}$ is the reference distance; $A_{0}$ is the intercept representing the path loss at the reference distance $d_{0}$; $n$ is the path loss exponent, which varies for different traffic scenarios and is set to $n=2.15$ for highway V2V scenarios; $X_{\sigma}$ represents shadow fading, modeled as a Gaussian random variable with zero mean and variance $\sigma^{2}$, caused by obstacles such as trees, buildings, and vehicles. Formally, $X_{\sigma}\sim\mathcal{N}(0,\sigma^{2})$, where the standard deviation is $\sigma=3$ dB.
The simulation environment models the communication scenario and control algorithms using Python, and integrates with SUMO for co-simulation and optimization via the TraCI interface. The simulation time step and duration are synchronized with those of SUMO. A summary of all simulation parameters is presented in Table I.
@@
During training, each batch consists of 60 time steps, and the maximum episode length is set to 600 steps. The learning rate for both the actor and critic networks is set to $1\times 10^{-3}$. All hyperparameters were determined through grid search.

### v2

@@ IV DCT-MARL Based Algorithm
In this section, the proposed DCT-MARL algorithm is described, as shown in Figure.1. The platoon control task is then formulated as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP). To ensure robust control performance under non-ideal communication conditions, the framework integrates two key mechanisms: (i) a dynamic communication topology that adaptively selects communication partners, and (ii) an explicit delay module that augments the state and reward design to account for time-varying transmission latency.
@@
The DCT-MARL algorithm is built upon the actor-critic architecture. Each actor consists of a multi-key gated communication network, which dynamically selects relevant neighbors based on inference mechanisms, and an action policy network that outputs control actions. At each control step, every agent gathers state information from vehicles within its communication range. Considering the possibility of communication failures, the agent adaptively determines its communication topology based on local observations, integrating self-state and received information for decision-making.
@@
The centralized critic aggregates joint actions and state representations from all agents to estimate each agent’s actionvalue function at each time step. An attention mechanism is utilized to extract relevant information from other agents’ observation-action pairs, enabling each agent to selectively filter critical environmental information. The critic network is trained via temporal difference learning [34] and is only used during training, enabling decentralized execution once the actors are deployed.
To optimize each agent $i$’s policy $\pi_{i}$, this paper minimizes the Kullback–Leibler (KL) divergence between the learned distribution $q_{\theta_{i}}$ and the target distribution $p$:
$-D_{\text{KL}}(q_{\theta_{i}}\parallel p)=\sum_{t}\mathbb{E}_{s_{t},u_{t},A_{t}\sim q_{\theta_{i}}}\Big{[}$ $r_{i}(s_{t},u_{t})+H\big{(}\pi_{i}(u_{i}\mid o_{i}(s,A))$
$\cdot\rho(A_{t}\mid s_{t})\big{)}\Big{]}.$ (8)
Beyond the reward term, the objective introduces an additional conditional entropy term for the joint distribution $\pi_{i}(a_{i}\mid o(s_{t},A_{t}))\,\rho(A_{t}\mid s_{t})$, which encourages exploration for agent $i$ and enhances the graph policy $\rho$. The adjacency trajectory matrices for the dynamic directed graph and its subgraphs are denoted as $\mathbf{A}$ and will be further detailed in the following section. The objective function for agent $i$ is thus redefined as:
@@
The critic network $Q_{i}$ is trained by minimizing the temporal-difference loss:
@@
During training, this paper follow these iterative procedures:
@@
where the binary label defined as: $y_{i}=\begin{cases}1,&\text{if }P_{\omega}(y_{i}\mid o_{i})\geq\delta\\ 0,&\text{otherwise},\end{cases}$ where $\delta$ is a tunable hyperparameter controlling the confidence threshold for classification. Notably, if a certain agent significantly impacts another agent’s state estimation during centralized training, its information must remain accessible during decentralized execution. To achieve this, the dynamic communication topology learned in training is retained for execution, enabling consistent and robust coordination.
@@ IV-A Dec-POMDP Formulation
State Space $S$: To mitigate the performance degradation caused by communication delays, this paper extend the definition of the agent’s state space by incorporating both historical control information and delay descriptors. Specifically, at time step $t$, the state of vehicle $i$ is defined as:
$S_{i,t}=\left\{p_{i,t},v_{i,t},acc_{i,t},u_{t-1},\xi_{i,t}\right\},$ (14)
where $p_{i,t}$, $v_{i,t}$, and $acc_{i,t}$ denote the vehicle’s current position, velocity, and acceleration, respectively; $u_{t-1}$ represents the most recent control input, and $\xi_{i,t}$ denotes the communication delay associated with the latest received information.
Although this formulation deviates from the conventional state definition, it is a rational extension under communication-impaired environments, where both the delayed control input and latency should be explicitly considered in the state representation. In connected vehicle systems, time-varying wireless communication often causes neighboring information to arrive with variable delays. If the control policy relies solely on such delayed observations, the Markov property may be compromised, thereby impairing decision-making performance.
@@
[…]
@@
$r_{i,k}=-\big{\{}$ $\omega_{1}\left(\text{err}_{p_{i},k}\right)+\omega_{2}\left(\text{err}_{v_{i},k}\right)+\omega_{3}\left(\text{jerk}_{i}\right)$ (15)
$+\omega_{4}\cdot\operatorname{ReLU}\left(d_{\text{safe}}-d_{i,k}\right)\big{\}},$ (16)
@@
$\text{jerk}_{i}=-\frac{\dot{acc}_{i}^{2}(t)}{\left[\frac{\left(u_{\max}-u_{\min}\right)}{\Delta t}\right]^{2}}.$ (17)
@@
In addition, considering the communication delay problem, each agent can only access the lagged observation information when making a decision, and cannot directly obtain the current true state. Therefore, this paper modify the reward function based on conditional expectation to adapt it to the optimal learning in the communication constrained environment. Specifically, for each agent $i$, in the extended state $S_{i,t}$, the reward function of delay is defined as follows:
$\tilde{R}_{i,t}\left(S_{i,t},acc_{i,k}^{\mathrm{CL}}\right)=\mathbb{E}_{x_{i,t}}\left[R_{i,t}\left(x_{i,t},acc_{i,t}\right)\mid S_{i,t}\right].$ (18)
@@
IV-B Dec-POMDP with Dynamic Communication topology
@@
Based on Dec-POMDP, this paper propose a communication topology using dynamic directed graphs. The structural properties of the graph are represented by the adjacency trajectory matrix $A$, which is denoted as $\mathbf{A}\in\mathbb{B}^{l\times n\times n}$, where B indicates the Boolean matrix, l denotes the length of the trajectory and n indicates the number of agents. If $\lceil A^{ij}_{t}\rceil=0$, then no communication occurred from agent $i$ to agent $j$ at time $t$. If $\lceil A^{ij}_{t}\rceil=1$, then communication transpired from agent $i$ to agent $j$ at time $t$.
IV-C Multi-Key Gating Network-Based Dynamic Communication topology
@@
During communication, each agent intuitively prefers to interact with those agents that significantly influence its decision-making strategy. Specifically, assuming agent $i$ can receive information from agent $j$, it first inputs its local observation $o_{i}$ along with the ID of agent $j$ into the model, which then outputs a confidence score indicating the likelihood of communication with agent $j$.
To quantify this causal influence, a feedforward neural network is employed to implement causal inference, leveraging the KL divergence to measure the causal confidence between agent $i$ and other agents:
$D_{\mathrm{KL}}\!\left(P(u_{i}\mid u_{-ij},o)\,\big{\|}\,P(u_{i},u_{j}\mid u_{-ij},o)\right),$ (20)
@@
where $P(u_{i}\mid u_{-ij},o)$ and $P(u_{i},u_{j}\mid u_{-ij},o)$ represent two conditional probability distributions over the action space of agent $i$, and $o$ denotes the joint observations of all agents. Here, $u_{-i}$ refers to the joint actions of all agents excluding agent $i$, and $u_{-ij}$ denotes the joint actions of all agents excluding both agents $i$ and $j$.
@@
The latter probability distribution does not depend on agent $j$’s actions, meaning that when making decisions, agent $i$ ignores the actions of agent $j$. Note that the conditional actions of other agents $u_{-i}$ are sampled from their current policies. The magnitude of KL divergence quantifies how agent $i$ would adjust its own policy if it considered agent $j$’s actions, thus indicating the degree of correlation between agent $j$’s policy and agent $i$’s policy.
@@
These probability distributions are computed via the joint action-value function. For instance, the softmax distribution over agent $i$’s actions is given by:
$P(u_{i}\mid u_{-ij},o)=\frac{\exp\!\left(Q(u_{i},u_{-ij},o)/\lambda\right)}{\sum_{u_{i}^{\prime}}\exp\!\left(Q(u_{i}^{\prime},u_{-ij},o)/\lambda\right)},$ (21)
@@
where $Q(u,o)$ is the joint action-value function, and $\lambda\in\mathbb{R}^{+}$ is the temperature parameter. The marginal distribution $P(u_{i}\mid u_{-ij},o)$ is computed as follows:
$P(u_{i}\mid u_{-ij},o)=\sum_{u_{j}}P(u_{i},u_{j}\mid u_{-ij},o).$ (22)
@@
After computing causal influence, a multi-key gating network is used to select $m$ communication agents. Using the Gumbel–Softmax activation, the network determines whether an agent communicates or not:
$\text{key}_{i}^{t}=\mathrm{gumbel\text{-}softmax}\!\left(W_{v}\tanh\!\left(W_{q}o_{i}^{t-1}+W_{k}o_{i}^{t-1}\right)\right),$ (23)
where Gumbel-Softmax [35, 36] is a binary activation (communicate or not), and the weight matrices $W_{v}$, $W_{q}$, and $W_{k}$ encode the communication relevance under dynamically changing platoon topologies.
@@
To improve interpretability and embed safety considerations, a distance weighting module assigns higher weights to closer vehicles, highlighting their critical role in collision avoidance:
$\omega_{j,t}=\frac{\frac{1}{2^{|i-j|}}}{\sum\limits_{l\in\mathcal{N}_{i,t}}\frac{1}{2^{|i-l|}}},$ (24)
where $d_{ij}$ is the distance between vehicles $i$ and $j$, and $\varepsilon$ is a small positive constant preventing division by zero.
Using the outputs of the gating network, the adjacency matrix $\mathbf{A}$ is then constructed by aggregating keys:
$A_{i,k}=\max\big{(}\text{key}_{i}^{t}\big{)},$ (25)
where $\text{key}_{i}^{k}$ denotes the number of keys assigned to agent $i$ at time step $t$, and $\lceil\cdot\rceil$ is the ceiling function. This process dynamically adjusts the communication topology, while the limited number of keys ensures that each agent only exchanges information with selected neighbors, keeping the input dimension stable. Finally, all received messages are passed through the message encoder network $e_{i}(\cdot)$ to obtain encoded messages $c_{i}$, which are fed into the policy network to generate the action distribution $\pi_{i}(u_{i}\mid c_{i},o_{i})$.
@@ V-A Experimental Setup
To validate the effectiveness of the proposed control strategy, a platoon of six vehicles is modeled within a 1-km highway segment in SUMO. The leading vehicle’s driving trajectory is derived from real-world traffic data collected on the eastbound I-80 highway in the San Francisco Bay Area, sourced from the NGSIM dataset [37]. As a widely recognized benchmark in platoon control research, the NGSIM dataset encompasses congested traffic scenarios, enabling a rigorous evaluation of the controller’s ability to manage stop-and-go traffic oscillations frequently encountered in real-world platoon operations. Integrating such realistic driving patterns facilitates more effective policy learning and enhances the applicability of the proposed approach to connected vehicle platoons operating under complex and dynamic conditions.
For the communication simulation, the wireless channel model incorporates both small-scale and large-scale fading effects. Small-scale fading is modeled as Rayleigh fading, which captures the rapid signal fluctuations caused by multipath propagation. Large-scale fading consists of path loss (PL) and shadow fading (SF), jointly modeling the attenuation of signal strength over distance and the slow variations caused by obstacles in the environment.
Path loss is modeled using the WINNER+ B1 model recommended by 3GPP [38], which is defined as:
$PL(d)=A_{0}+10n\log_{10}\!\left(\frac{d}{d_{0}}\right)+X_{\sigma},\quad\text{(dB)}$ (26)
where $d_{0}$ is the reference distance; $A_{0}$ is the intercept representing the path loss at the reference distance $d_{0}$; $n$ is the path loss exponent, which varies for different traffic scenarios and is set to $n=2.15$ for highway V2V scenarios; $X_{\sigma}$ represents shadow fading, modeled as a Gaussian random variable with zero mean and variance $\sigma^{2}$, caused by obstacles such as trees, buildings, and vehicles. Formally,
$X_{\sigma}\sim\mathcal{N}(0,\sigma^{2}),\qquad\sigma=3\ \text{dB}.$ (27)
The simulation environment models the communication scenario and control algorithms using Python, and integrates with SUMO for co-simulation and optimization via the TraCI interface. The simulation time step and total duration are synchronized with SUMO to ensure consistency. Table LABEL:tab:sim_params summarizes the key simulation parameters.
@@
Highway Length 1000m
Minimum Vehicle Velocity 0 m/s
Maximum Vehicle Velocity 30 m/s
Acceleration Range -3 m/s2 to 3 m/s2
Platoon Vehicle Number 6
@@
Training is conducted for 1 million steps in a simulated traffic environment, with an episode length of $600$ steps. The discount factor is set to $\gamma=0.99$, and the learning rates for the actor and critic networks are $3.0\times 10^{-4}$ and $2.5\times 10^{-4}$, respectively. Each algorithm is trained three times using different random seeds to ensure robustness and generalization, and all hyperparameters are selected through a grid search process.